# Towards Understanding Why Label Smoothing Degrades Selective Classification and How to Fix It

### v2

@@ Abstract
Label smoothing (LS) is a popular regularisation method for training neural networks as it is effective in improving test accuracy and is simple to implement. “Hard” one-hot labels are “smoothed” by uniformly distributing probability mass to other classes, reducing overfitting. Prior work has suggested that in some cases LS can degrade selective classification (SC) – where the aim is to reject misclassifications using a model’s uncertainty. In this work, we first demonstrate empirically across an extended range of large-scale tasks and architectures that LS consistently degrades SC. We then address a gap in existing knowledge, providing an explanation for this behaviour by analysing logit-level gradients: LS degrades the uncertainty rank ordering of correct vs incorrect predictions by regularising the max logit more when a prediction is likely to be correct, and less when it is likely to be wrong. This elucidates previously reported experimental results where strong classifiers underperform in SC. We then demonstrate the empirical effectiveness of post-hoc logit normalisation for recovering lost SC performance caused by LS. Furthermore, linking back to our gradient analysis, we again provide an explanation for why such normalisation is effective.
@@ 1 Introduction
%coverage@risk=1$\uparrow$ 17.69 0.2 0.53 0.49 21.9
@@
    We show empirically, across a range of large-scale architectures (CNN,ViT) and tasks (image classification, semantic segmentation), that training with LS consistently leads to degraded SC performance (see Fig. 1), even if it may improve accuracy. Moreover, we show that the degradation worsens with stronger LS. As LS can be found in the training recipes of many of the models evaluated in (Galil et al., 2023; Cattelan & Silva, 2024), this suggests LS as one potential cause for previously unexplained negative results where strong classifiers underperform at SC.
@@
    We address a gap in the understanding of LS by providing an explanation of this behaviour through analysing the logit-level gradients of the LS loss. We show that the amount LS regularises/suppresses the max logit directly corresponds to the true probability of error $P_{\text{error}}$, with regularisation increasing(decreasing) the more likely a prediction is correct(wrong). This leads to relatively higher uncertainty on correct predictions and lower uncertainty on misclassifications, degrading the ranking of uncertainties and hurting SC compared to vanilla CE.
@@
    We show that post-hoc logit normalisation (Cattelan & Silva, 2024) is effective in negating the degradation from LS (Fig. 1). Moreover, we elucidate this effectiveness through the lens of our gradient-based analysis. Linking back to the imbalanced logit regularisation of LS, we find that logit normalisation compensates for this effect by increasing uncertainty as the max logit increases.
@@ 2 Preliminaries
Consider a $K$-class neural network classifier with parameters $\boldsymbol{\theta}$ that models the conditional distribution $P(y|\boldsymbol{x};\boldsymbol{\theta})$ of labels $y\in\mathcal{Y}=\{\omega_{k}\}_{k=1}^{K}$ given inputs $\boldsymbol{x}\in\mathcal{X}=\mathbb{R}^{D}$. Typically the network has a categorical softmax output $\boldsymbol{\pi}(\boldsymbol{x};\boldsymbol{\theta})\in[0,1]^{K}$,
@@
where $\delta$ is the Kronecker delta and KL[||][$\cdot$|$\cdot$][ ⋅ | | ⋅ ] is the Kullback–Leibler divergence. We use $\bar{\boldsymbol{\pi}}(\boldsymbol{x})\in[0,1]^{K}$ as a shorthand for the true conditional categorical, i.e. $\bar{\pi}_{k}=P_{\text{data}}(\omega_{k}|\boldsymbol{x})$. Predictions $\hat{y}$ are then made on new unlabelled input data $\boldsymbol{x}^{*}$ using classifier function $f$ during deployment,
@@ 2.2 Selective classification (SC)
i.e. the (negative of the) Maximum Softmax Probability (MSP) (Hendrycks & Gimpel, 2017) is the model’s estimate of the probability of its prediction $\hat{y}$. We focus on this uncertainty measure $U$ for SC as it is a popular default choice and has been shown to consistently and reliably perform well in the literature (Xia & Bouganis, 2022b; 2023; Jaeger et al., 2023; Feng et al., 2023) compared to alternatives. MSP is a natural choice as $P(y|\boldsymbol{x};\boldsymbol{\theta})$ models $P_{\text{data}}(y|\boldsymbol{x})$, and $U=-\max_{k}P_{\text{data}}(\omega_{k}|\boldsymbol{x})$ (or any scalar monotonic to it) provides the optimal risk if the true distribution $P_{\text{data}}(y|\boldsymbol{x})$ is known (Chow, 1970). We omit evaluations on other softmax-based $U$ from the main paper as we find them to behave similarly to MSP. In line with previous work (Xia & Bouganis, 2022b; Jaeger et al., 2023; Zhu et al., 2024), we find that OOD detection (Yang et al., 2021) scores perform poorly at SC, and also omit them (see Sec. E.1 for discussion and additional results).
@@ 3 The effect of label smoothing on selective classification
Experimental details. We investigate large-scale [^2] We additionally provide $32\times 32$ small-scale CIFAR experiments in Sec. A.3. image classification on ImageNet-1k (Russakovsky et al., 2015) and semantic segmentation (pixel-level classification) on Cityscapes (Cordts et al., 2016). For ImageNet, we evaluate on the original validation set and randomly split 50,000 images from the training set for validation. For Cityscapes, we randomly split the original validation set into 100 validation and 400 evaluation images. To estimate the risk, we extract the model output prior to the final interpolation and subsample 5000 labelled pixels per image at random. We evaluate on the same pixels between models. The only parameter varied between training runs is the level of LS $\alpha$, and to isolate the effects of LS, we train all models from scratch using simple recipes. We purposely avoid augmentations such as MixUp (Zhang et al., 2018) and CutMix (Yun et al., 2019) as these directly affect the training labels, which would interfere with our experiments. For ImageNet classification, we train ResNet-50 (He et al., 2016) and ViT-S-16 (Dosovitskiy et al., 2021) using only random resized cropping and flipping for data augmentation. To achieve decent accuracy for ViT training from scratch without advanced augmentations, we use sharpness-aware minimisation (SAM) (Foret et al., 2021; Chen et al., 2022). For semantic segmentation on Cityscapes, we train DeepLabV3+ (Chen et al., 2018) (ResNet-101 backbone) using only random cropping, flipping and colour jitter for augmentations. Full training details for reproducibility can be found in the Appendix B. We will release our code in the future.
@@ 3.1 Label smoothing degrades selective classification
Figure 3: Risk-coverage plots for different levels of LS $\alpha$ for different models and tasks (ImageNet classification and Cityscapes semantic segmentation). Although LS may improve error rate at 100% coverage, increasing $\alpha$ increasingly and consistently degrades SC performance.
@@
Upon inspection, many of the SC-underperforming models benchmarked in (Galil et al., 2023; Cattelan & Silva, 2024) are trained using LS, aligning with our results. The models are sourced from repositories such as torchvision (Paszke et al., 2019) and timm (Wightman, 2019), where the training recipes are optimised for top-1 accuracy. We note that for ImageNet, LS is such a common technique that it is often used by default, and not even mentioned in papers, e.g. (Tan & Le, 2019; 2021). Overall, our results highlight that only optimising for accuracy may result in negative downstream consequences and that practitioners of SC need to be aware of the effects of their training recipes. We include additional discussion of existing benchmarks and model checkpoints in Sec. E.4.
@@ 4 Towards explaining the negative effect of LS on SC
In the following section, we attempt to delve deeper into why our empirical results (and those in (Zhu et al., 2024)) occur, shedding light on a previously unexplained phenomenon. It might intuitively seem odd that LS degrades SC performance as the transform applied to the true targets in Eq. 8 $(1-\alpha)\bar{\boldsymbol{\pi}}(\boldsymbol{x})+\alpha\boldsymbol{u}$ reduces the max probability for all samples but does not change the relative ranking. However, this only carries over to the model when it is knowledgeable about the data, so it is well fit $\boldsymbol{\pi}(\boldsymbol{x};\boldsymbol{\theta})\approx(1-\alpha)\bar{\boldsymbol{\pi}}(\boldsymbol{x})+\alpha\boldsymbol{u}$. For such samples, the model has learnt to be more uncertain. However, we cannot assume that the model is equally well fit on all regions of the data distribution. [^3] For the sake of simplicity we avoid the term epistemic uncertainty (see Sec. E.2 for brief discussion).
@@ 4.1 Comparing logit gradients between CE and LS
To better understand how LS could lead to degraded SC, we consider how LS affects logit-level training gradients. These are the first term in the chain rule for backpropagation and so directly contribute to all parameter gradients during training. We take the gradient of $\mathcal{L}^{\text{true}}$ (Eqs. 4 and 8), [^4] Here we assume that the empirical loss $\mathcal{L}$ (Eqs. 2 and 6) approximates $\mathcal{L}^{\text{true}}$, in order to relate our discussion to $P_{\text{error}}$. However, the same analysis on the empirical loss leads to similar conclusions (see Appendix C).
@@
for a single sample, where in a slight abuse of notation we omit the outer expectation over $p_{\text{data}}(\boldsymbol{x})$ for convenience. We can then define the regularisation gradient on the logits,
$\frac{\partial\mathcal{L}^{\text{true}}_{\text{reg}}}{\partial v_{k}}=\frac{\partial(\mathcal{L}_{\text{LS}}^{\text{true}}-\mathcal{L}_{\text{CE}}^{\text{true}})}{\partial v_{k}}=\frac{\mathcal{L}_{\text{LS}}^{\text{true}}}{\partial v_{k}}-\frac{\mathcal{L}_{\text{CE}}^{\text{true}}}{\partial v_{k}}=\alpha\left[\bar{\pi}_{k}-1/K\right]=\alpha\bar{\pi}_{k}-\alpha/K~{},$ (14)
@@
4.2 Imbalanced regularisation degrades uncertainty ranking of  vs
@@
Figure 4: How the regularisation gradient (Eq. 14), which is the difference between LS and CE gradients, affects the logits differently. LS affects the max logit differently depending on how well fit the model is for a given sample $\boldsymbol{x}$. On the left when $U$ is lower (sharper softmax), the regularisation on the max logit is lower when the model is poorly fit and likely to be wrong. On the right when $U$ is higher (flatter softmax), the regularisation is higher when the model is poorly fit and more likely to be correct. Thus, LS degrades the softmax’s ability to separate  vs  , hurting SC.
We now consider how the regularisation gradient affects the maximum logit $v_{\text{max}}$,
@@
$\frac{\partial\mathcal{L}^{\text{true}}_{\text{reg}}}{\partial v_{\text{max}}}=\alpha\bar{\pi}_{\hat{y}}-\alpha/K=\alpha[1-P_{\text{error}}]-\alpha/K,$ (15)
@@
which shows that the regularisation on $v_{\text{max}}$ decreases as the probability of error increases. This directly impacts softmax-based $U$ such as MSP (see Eq. 1), especially for lower uncertainties where $v_{\text{max}}$ dominates the softmax. Eq. 15 suggests that for predictions that share the same value of $U$, those with higher $P_{\text{error}}$ (more likely ✗), will have $v_{\text{max}}$ regularised less, whilst predictions with lower $P_{\text{error}}$ (more likely ✓) will have $v_{\text{max}}$ regularised more. Thus, softmax-based $U$ will have the relative ranking of correct ✓ and incorrect ✗ predictions degraded – less uncertain on ✗ , more uncertain on ✓. To further build an intuition, we consider how a model may not be uniformly well-fit over the data distribution during training. As illustrated in Fig. 4, varying levels of fit with regards to the true distribution $\bar{\pi}$ leads to differing $P_{\text{error}}$ for predictions with similar $U$, resulting in the degradation described above. Finally, as the regularisation gradient in Eq. 15 is proportional to $\alpha$, this explains why stronger LS leads to greater degradation. In summary: for a given predicted $U$ the $P_{\text{error}}$ will vary depending on how well fit the model is. LS suppresses the max logit more(less) when $P_{\text{error}}$ is lower(higher), degrading the ranking of  vs  for softmax-based $U$, hurting SC. We note that although our analysis is performed on training gradients, all our experiments are performed on evaluation data, suggesting that effects indeed generalise to unseen data. We also note that our explanation is purely based on the loss, and thus generalises across network architectures and tasks, matching the empirical results in Sec. 3.
@@
LS empirically leads to higher $v_{\text{max}}$ on misclassifications . To further validate the effects of Eq. 15, we plot the mean±$\pm$±std. of $v_{\text{max}}$ given $\pi_{\text{max}}$ for ResNet-50 on evaluation data in Fig. 5. We see that for LS the distribution of correct ✓  is below incorrect ✗, whilst for CE they are roughly similar. This provides further empirical support for Eq. 15 which states that when using LS, $v_{\text{max}}$ is more strongly suppressed for lower $P_{\text{error}}$. For results on other models see Sec. A.2.
@@
LS empirically leads to increasing overconfidence on less-well-fit data. To further investigate how LS affects uncertainty differently depending on how well-fit/knowledgeable the model is about data, we artificially introduce distribution-shifted data that we expect our models to be worse fit on. We use ImageNet-Sketch (Wang et al., 2019), a dataset containing sketches of each ImageNet class and evaluate on the combination of the 50,889 ImageNet-Sketch and 50,000 ImageNet evaluation images in Fig. 6. Even though the regularisation of LS improves the error rate, the degradation of SC at lower coverages from LS is exacerbated by the distribution-shifted data. At 10% coverage of the combined evaluation set, LS leads to accepting many more errors compared to CE, with an increasing proportion originating from ImageNet-Sketch. That is to say, LS leads to increasing overconfidence on less-well-fit data (ImageNet-Sketch). This empirically highlights the situation illustrated on the left of Fig. 4 where confident and well-fit samples have their max logit suppressed, whilst confident but poorly fit samples do not, leading to samples with higher $P_{\text{error}}$ being relatively less uncertain (overconfident). For results on other models see Sec. A.2.
@@
Figure 5: Distribution of the max logit $v_{\text{max}}$ given the MSP $\pi_{\text{max}}$ for correct  and incorrect  predictions on evaluation data. $v_{\text{max}}$ is lower for  for the LS model, whilst the distributions are roughly similar for CE. This empirically matches the imbalanced max logit regularisation described in Eq. 15. We calculate the mean±$\pm$±std. in a 0.05-wide sliding window.
@@
CE #samples 8994 1094
#errors 93 80
error rate 1.0 7.3
LS $\alpha=0.1$samples 8561 1527
#errors 187 211
error rate 2.2 13.8
LS $\alpha=0.2$samples 8171 1917
#errors 244 365
error rate 3.0 19.0
LS $\alpha=0.3$samples 7915 2173
#errors 297 541
error rate 3.8 24.9

### v3

@@ Abstract
Label smoothing (LS) is a popular regularisation method for training neural networks as it is effective in improving test accuracy and is simple to implement. “Hard” one-hot labels are “smoothed” by uniformly distributing probability mass to other classes, reducing overfitting. Prior work has suggested that in some cases LS can degrade selective classification (SC) – where the aim is to reject misclassifications using a model’s uncertainty. In this work, we first demonstrate empirically across an extended range of large-scale tasks and architectures that LS consistently degrades SC. We then address a gap in existing knowledge, providing an explanation for this behaviour by analysing logit-level gradients: LS degrades the uncertainty rank ordering of correct vs incorrect predictions by suppressing the max logit more when a prediction is likely to be correct, and less when it is likely to be wrong. This elucidates previously reported experimental results where strong classifiers underperform in SC. We then demonstrate the empirical effectiveness of post-hoc logit normalisation for recovering lost SC performance caused by LS. Furthermore, linking back to our gradient analysis, we again provide an explanation for why such normalisation is effective. Project page: https://ensta-u2is-ai.github.io/Understanding-Label-smoothing-Selective-classification/
@@ 1 Introduction
%coverage@risk=1$\uparrow$ 15.66 0.05 0.06 0.02 22.31
@@
    We show empirically, across a range of large-scale architectures (CNN,ViT) and tasks (image classification, semantic segmentation), that training with LS consistently leads to degraded SC performance (see Fig. 1), even if it may improve accuracy. Moreover, we find that the degradation worsens with stronger LS. As LS can be found in the training recipes of many of the models evaluated in (Galil et al., 2023; Cattelan & Silva, 2024), this suggests LS as one potential cause for previously unexplained negative results where strong classifiers underperform at SC.
@@
    We address a gap in the understanding of LS by providing an explanation of this behaviour through analysing the logit-level gradients of the LS loss. We show that the amount LS suppresses the max logit directly corresponds to the true probability of error $P_{\text{error}}$, with suppression increasing(decreasing) the more likely a prediction is correct(wrong). This leads to relatively higher uncertainty on correct predictions and lower uncertainty on misclassifications, degrading the ranking of uncertainties and hurting SC compared to vanilla CE.
@@
    We show that post-hoc logit normalisation (Cattelan & Silva, 2024) is effective in negating the degradation from LS (Fig. 1). Moreover, we elucidate this effectiveness through the lens of our gradient-based analysis. Linking back to the imbalanced logit suppression of LS, we find that logit normalisation compensates for this effect by increasing uncertainty as the max logit increases.
@@ 2 Preliminaries
For a glossary of notation see Appendix A. Consider a $K$-class neural network classifier with parameters $\boldsymbol{\theta}$ that models the conditional distribution $P(y|\boldsymbol{x};\boldsymbol{\theta})$ of labels $y\in\mathcal{Y}=\{\omega_{k}\}_{k=1}^{K}$ given inputs $\boldsymbol{x}\in\mathcal{X}=\mathbb{R}^{D}$. Typically the network has a categorical softmax output $\boldsymbol{\pi}(\boldsymbol{x};\boldsymbol{\theta})\in[0,1]^{K}$,
@@
where $\delta_{ij}=1\text{ if }i=j,\text{ and }0\text{ if }i\neq j$ is the Kronecker delta and KL[||][$\cdot$|$\cdot$][ ⋅ | | ⋅ ] is the Kullback–Leibler divergence. We use $\bar{\boldsymbol{\pi}}(\boldsymbol{x})\in[0,1]^{K}$ as a shorthand for the true conditional categorical, i.e. $\bar{\pi}_{k}=P_{\text{data}}(\omega_{k}|\boldsymbol{x})$. Predictions $\hat{y}$ are then made on new unlabelled input data $\boldsymbol{x}^{*}$ using classifier function $f$ during deployment,
@@ 2.2 Selective classification (SC)
i.e. the (negative of the) Maximum Softmax Probability (MSP) (Hendrycks & Gimpel, 2017) is the model’s estimate of the probability of its prediction $\hat{y}$. We focus on this uncertainty measure $U$ for SC as it is a popular default choice and has been shown to consistently and reliably perform well in the literature (Xia & Bouganis, 2022b; 2023; Jaeger et al., 2023; Feng et al., 2023) compared to alternatives. MSP is a natural choice as $P(y|\boldsymbol{x};\boldsymbol{\theta})$ models $P_{\text{data}}(y|\boldsymbol{x})$, and $U=-\max_{k}P_{\text{data}}(\omega_{k}|\boldsymbol{x})$ (or any scalar monotonic to it) provides the optimal risk if the true distribution $P_{\text{data}}(y|\boldsymbol{x})$ is known (Chow, 1970). We omit evaluations on other softmax-based $U$ from the main paper as we find them to behave similarly to MSP. In line with previous work (Xia & Bouganis, 2022b; Jaeger et al., 2023; Zhu et al., 2024), we find that OOD detection (Yang et al., 2021) scores perform poorly at SC, and also omit them (see Sec. F.1 for discussion and additional results).
@@ 3 The effect of label smoothing on selective classification
Experimental details. We investigate large-scale [^2] We additionally provide small-scale CIFAR, tabular and text experiments in Secs. B.3, B.4 and B.5. image classification on ImageNet-1k (Russakovsky et al., 2015) and semantic segmentation (pixel-level classification) on Cityscapes (Cordts et al., 2016). For ImageNet, we randomly split the validation set into 10,000 validation and 40,000 evaluation images. For Cityscapes, we randomly split the original validation set into 100 validation and 400 evaluation images. To estimate the risk for the RC curves, we subsample 5000 labelled pixels per image at random. We evaluate on the same pixels between models. The only parameter varied between training runs is the level of LS $\alpha$, and to isolate the effects of LS, we train all models from scratch using simple recipes. We purposely avoid augmentations such as MixUp (Zhang et al., 2018) and CutMix (Yun et al., 2019) as these directly affect the training labels, which would interfere with our experiments. For ImageNet classification, we train ResNet-50 (He et al., 2016) and ViT-S-16 (Dosovitskiy et al., 2021) using only random resized cropping and flipping for data augmentation. To achieve decent accuracy for ViT training from scratch without advanced augmentations, we use sharpness-aware minimisation (SAM) (Foret et al., 2021; Chen et al., 2022). For semantic segmentation on Cityscapes, we train DeepLabV3+ (Chen et al., 2018) (ResNet-101 backbone) using only random cropping, flipping and colour jitter for augmentations. Extended training details for reproducibility can be found in the Appendix C. We release our code on GitHub.
@@ 3.1 Label smoothing degrades selective classification
Figure 3: Risk-coverage plots for different levels of LS $\alpha$ for different models and tasks (ImageNet classification and Cityscapes semantic segmentation). Although it may improve error rate/accuracy at 100% coverage, label smoothing consistently degrades SC performance.
@@
Upon inspection, many of the SC-underperforming models benchmarked in (Galil et al., 2023; Cattelan & Silva, 2024) are trained using LS, aligning with our results. The models are sourced from repositories such as torchvision (Paszke et al., 2019) and timm (Wightman, 2019), where the training recipes are optimised for top-1 accuracy. We note that for ImageNet, LS is such a common technique that it is often used by default, and not even mentioned in papers, e.g. (Tan & Le, 2019; 2021). Overall, our results highlight that only optimising for accuracy may result in negative downstream consequences and that practitioners of SC need to be aware of the effects of their training recipes. We include additional discussion of existing benchmarks and model checkpoints in Sec. F.6.
@@ 4 Towards explaining the negative effect of LS on SC
In the following section, we attempt to delve deeper into why our empirical results (and those in (Zhu et al., 2024)) occur, shedding light on a previously unexplained phenomenon. It might intuitively seem odd that LS degrades SC performance as the transform applied to the true targets in Eq. 8 $(1-\alpha)\bar{\boldsymbol{\pi}}(\boldsymbol{x})+\alpha\boldsymbol{u}$ reduces the max probability for all samples but does not change the relative ranking. However, this only carries over to the model when it is knowledgeable about the data, so it is well fit $\boldsymbol{\pi}(\boldsymbol{x};\boldsymbol{\theta})\approx(1-\alpha)\bar{\boldsymbol{\pi}}(\boldsymbol{x})+\alpha\boldsymbol{u}$. For such samples, the model has learnt to be more uncertain. However, we cannot assume that the model is equally well fit on all regions of the data distribution. [^3] For the sake of simplicity we avoid the term epistemic uncertainty (see Sec. F.2 for brief discussion).
@@ 4.1 Comparing logit gradients between CE and LS
To better understand how LS could lead to degraded SC, we consider how LS affects logit-level training gradients. These are the first term in the chain rule for backpropagation and so directly contribute to all parameter gradients during training. We take the gradient of $\mathcal{L}^{\text{true}}$ (Eqs. 4 and 8), [^4] Here we assume that the empirical loss $\mathcal{L}$ (Eqs. 2 and 6) approximates $\mathcal{L}^{\text{true}}$, in order to relate our discussion to $P_{\text{error}}$. However, the same analysis on the empirical loss leads to similar conclusions (see Appendix D).
@@
for a single sample, where in a slight abuse of notation we omit the outer expectation over $p_{\text{data}}(\boldsymbol{x})$ for convenience. We can then define the suppression gradient on the logits,
$\frac{\partial\mathcal{L}^{\text{true}}_{\text{{\color[rgb]{0,0,0}sup}}}}{\partial v_{k}}=\frac{\partial(\mathcal{L}_{\text{LS}}^{\text{true}}-\mathcal{L}_{\text{CE}}^{\text{true}})}{\partial v_{k}}=\frac{\partial\mathcal{L}_{\text{LS}}^{\text{true}}}{\partial v_{k}}-\frac{\partial\mathcal{L}_{\text{CE}}^{\text{true}}}{\partial v_{k}}=\alpha\left[\bar{\pi}_{k}-1/K\right]=\alpha\bar{\pi}_{k}-\alpha/K~{},$ (14)
@@
4.2 Imbalanced suppression degrades uncertainty ranking of  vs
@@
Figure 4: How the suppression gradient (Eq. 14), i.e. the difference between LS and CE gradients, affects the logits differently. LS affects the max logit differently depending on how well fit the model is for a given sample $\boldsymbol{x}$. In the left two when $U$ is lower (sharper softmax), the suppression on the max logit is lower when the model is poorly fit and likely to be wrong. In the right two when $U$ is higher (flatter softmax), the suppression is higher when the model is poorly fit and more likely to be correct. Thus, LS degrades the softmax’s ability to separate  vs  , hurting SC.
We now consider how the suppression gradient affects the maximum logit $v_{\text{max}}$,
@@
$\frac{\partial\mathcal{L}^{\text{true}}_{\text{{\color[rgb]{0,0,0}sup}}}}{\partial v_{\text{max}}}=\alpha\bar{\pi}_{\hat{y}}-\alpha/K=\alpha[1-P_{\text{error}}]-\alpha/K,$ (15)
which shows that the suppression on $v_{\text{max}}$ decreases as the probability of error increases. This directly impacts softmax-based $U$ such as MSP (see Eqs. 1 and 12), as the exponentiation of the softmax will be dominated by the largest logits, especially the max logit $v_{\text{max}}$. The max logit will be especially dominant on high-confidence samples which are the last to be rejected.
@@
Eq. 15 suggests that for predictions that share the same value of $U$, those with higher $P_{\text{error}}$ (more likely ✗), will have $v_{\text{max}}$ suppressed less, whilst predictions with lower $P_{\text{error}}$ (more likely ✓) will have $v_{\text{max}}$ suppressed more. Thus, softmax-based $U$ will have the relative ranking of correct ✓ and incorrect ✗ predictions degraded – less uncertain on ✗ , more uncertain on ✓. To further build an intuition, we consider how a model may not be uniformly well-fit over the data distribution during training. As illustrated in Fig. 4, varying levels of fit with regards to the true distribution $\bar{\pi}$ leads to differing $P_{\text{error}}$ for predictions with similar $U$, resulting in the degradation described above. Finally, as the suppression gradient in Eq. 15 is proportional to $\alpha$, this explains why stronger LS leads to greater degradation. In summary: for a given predicted $U$ the $P_{\text{error}}$ will vary depending on how well fit the model is. LS suppresses the max logit more(less) when $P_{\text{error}}$ is lower(higher), degrading the ranking of  vs  for softmax-based $U$, hurting SC. We note that although our analysis is performed on training gradients, all our experiments are performed on evaluation data, suggesting that effects indeed generalise to unseen data. We also note that our explanation is purely based on the loss, and thus generalises across network architectures and tasks, matching the empirical results in Sec. 3.
@@
LS empirically leads to higher $v_{\text{max}}$ on misclassifications . To further validate the effects of Eq. 15, we plot the mean±$\pm$±std. of $v_{\text{max}}$ given $\pi_{\text{max}}$ for ResNet-50 on evaluation data in Fig. 5. We see that for LS the distribution of correct ✓  is below incorrect ✗, whilst for CE they are roughly similar. This provides further empirical support for Eq. 15 which states that when using LS, $v_{\text{max}}$ is more strongly suppressed for lower $P_{\text{error}}$. For results on other models see Sec. B.2.
@@
LS empirically leads to increasing overconfidence on less-well-fit data. To further investigate how LS affects uncertainty differently depending on how well-fit/knowledgeable the model is about data, we artificially introduce distribution-shifted data that we expect our models to be worse fit on. We use ImageNet-Sketch (Wang et al., 2019), a dataset containing sketches of each ImageNet class and evaluate on the combination of the 50,889 ImageNet-Sketch and 40,000 ImageNet evaluation images in Fig. 6. Even though the regularisation of LS improves the error rate, the degradation of SC at lower coverages from LS is exacerbated by the distribution-shifted data. At 10% coverage of the combined evaluation set, LS leads to accepting many more errors compared to CE, with an increasing proportion originating from ImageNet-Sketch. That is to say, LS leads to increasing overconfidence on less-well-fit data (ImageNet-Sketch). This empirically highlights the situation illustrated on the left of Fig. 4 where confident and well-fit samples have their max logit suppressed, whilst confident but poorly fit samples do not, leading to samples with higher $P_{\text{error}}$ being relatively less uncertain (overconfident). For results on other models see Sec. B.2.
@@
Figure 5: Distribution of the max logit $v_{\text{max}}$ given the MSP $\pi_{\text{max}}$ for correct  and incorrect  predictions on evaluation data. $v_{\text{max}}$ is lower for  for the LS model, whilst the distributions are roughly similar for CE. This empirically matches the imbalanced max logit suppression described in Eq. 15. We calculate the mean±$\pm$±std. in a 0.05-wide sliding window.
@@
CE #samples 7940 1148
#errors 94 74
error rate 1.2 6.4
LS $\alpha=0.1$samples 7422 1666
#errors 174 296
error rate 2.3 17.8
LS $\alpha=0.2$samples 6808 2280
#errors 225 549
error rate 3.3 24.1
LS $\alpha=0.3$samples 6726 2362
#errors 258 647
error rate 3.8 27.4